Status Report

NASA ISS On-Orbit Status 24 October 2009

By SpaceRef Editor
October 23, 2009
Filed under , , ,
NASA ISS On-Orbit Status 24 October 2009

All ISS systems continue to function nominally, except those noted previously or below. Saturday – crew light-duty day.

At day begin, FE-1 Suraev did the regular daily check of the new aerosol filters at the Russian Elektron O2 generator which he installed on 10/19 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) and hydrogen outlet pipe (filter FA-V). [Maxim again inspects the filters tonight at bedtime, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]

FE-2 Stott removed the alignment guides from the CIR (Combustion Integrated Rack) in the Lab to allow PaRIS (Passive Rack Isolation System) activation by the ground for operations requiring a microgravity environment. Tonight before bedtime, Nicole will re-install the guides to lock PaRIS down again.

The six-member crew performed the regular weekly three-hour task of thorough station cleaning, including COL (Columbus Orbital Laboratory) and Kibo. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the sleep stations with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

FE-1 Suraev did the periodic checkup behind ASU panel 139 in the SM (Service Module) on a fluid connector (MNR-NS) of the SM-U urine collection system, looking for potential moisture.

As part of the house cleaning, Suraev & FE-3 Romanenko conducted regular maintenance inspection & cleaning on fan screens, Group A, in the FGB (TsV2), DC1 (V3), and SM (VPkhO, VPrK, FS5, FS6 & FS9).

Later, Maxim also performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP-Moscow. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Roman completed the regular maintenance of the Russian POTOK-150MK (150 micron) air purification subsystems in the SM and FGB by cleaning the pre-filters with a vacuum cleaner with narrow nozzle attachment and later restarting POTOK in automatic mode.

FE-4 Thirsk completed the now regular transfer of pre-treated Russian urine directly from EDV-U container into the Lab UPA (Urine Processor Assembly) for processing, done each morning for about 5 minutes.

Before undertaking the transfer, Thirsk moved the UPA/EDV-U pump unit’s power cable from its plug-in in Locker 2 of ER-6 (EXPRESS Rack 6) to the ER-6 Locker 4 power outlet. [Locker 2 is a 10 amp power source, but the 22-gauge EDV-U cable is rated for a lower current limit and has no built-in fuse to protect against overcurrent in the event of a short circuit inside the compressor. The Locker 6 outlet has been set via software to 5 amp power outlet. This cable will be replaced with a new design on Mission 20A when the system is moved to Node 3.]

For today’s several VolSci (Voluntary Weekend Science) activities, on the crew’s free time, Nicole Stott & Frank De Winne set up the hardware for another BISE (Bodies in the Space Environment) experiment run, then each of them worked through the protocol and taking photographs of the partner. [The CSA (Canadian Space Agency)-sponsored BISE experiment studies how astronauts perceive Up and Down in microgravity, investigating the relative contributions of internal & external cues to self-orientation during and after micro-G exposure. BISE data collection must be performed at least one hour after any exercise. The specific objective of the BISE project is to conduct experiments during long-duration micro-G conditions to better understand how humans first adapt to micro-G and then re-adapt to normal gravity conditions upon return to earth. This experiment involves comparisons of preflight, flight, and post-flight perceptions and mental imagery, with special reference to spaceflight-related decreases in the vertical component of percepts. The test involves having subjects view a computer screen through a cylinder that blocks all other visual information. The astronauts are being presented with background images with different orientations relative to their bodies.]

For a second VolSci activity, Bob Thirsk set up the G1 camcorder for recording his subsequent performance of an educational science Podcast script describing ESA’s SODI-IVIDIL (Selectable Optical Diagnostics Instrument/Influence of Vibration on Diffusion in Liquids) and Canada’s MVIS (Microgravity Vibration Insulation System) activities. [Without buoyancy in the station’s weightless environment, molecules can be studied in new ways, isolating processes that may have been obscured in gravity on Earth. SODI-IVIDIL uses two motive forces – vibration & thermo-diffusion – to learn more about molecular movement. To achieve that, inside the MSG (Microgravity Science Glovebox) liquid samples are being shaken and also heated & cooled. With SODI, a sophisticated optical instrument, temperature and concentration of the molecules are then being measured. Except for purposes like SODI-IVIDIL, vibrations (hum of the ventilation system, pounding of the exercise machines, grinding of revolving solar panels) are unwanted on the ISS. To counter this problem, Canada has developed a technology called MVIS. It uses magnetic actuators to levitate experiments, held in FSL (Fluid Science Laboratory), so that they are not affected by vibrations. Sensors and an onboard computer control the position of the contained experiments, ensuring they remain free-floating. Podcast: a series of digital media files (either audio or video) that are released episodically and downloaded through web syndication or on demand.]

Bob also had ~30 min set aside to review material related to BCAT (Binary Colloidal Alloy Test) operations and procedures for the planned setting up of the necessary equipment in the Kibo JPM (JEM Pressurized Module). [Earlier this year (May) the crew deemed the US Lab too crowded for running BCAT. Starting 10/26, BCAT will be running in the JPM on the forward side of the airlock.]

Afterwards, Thirsk shut down the EK (EarthKAM) system in Node-1 and stowed the equipment, then relocated the Ku-band power supply used for EK from Node-1 to the JPM for powering BCAT beginning 10/26. [EK successfully conducted 5-days of image gathering for 122 Education Groups (Elementary through Collegiate levels) for 9761 students. 97 of the groups are from US Schools, 10 from Australia, the remainder from Spain: 2, Germany: 2, Japan, Dominican Republic, France, Lebanon, United Kingdom, Finland, India, Denmark, Netherlands, Romania, Guatemala. Over 750 images have been downlinked with two more days of operations still to follow.]

Nicole, Bob, Jeff & Frank filled out their regular weekly FFQs (Food Frequency Questionnaires) on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]

Nicole also completed the regular bi-monthly reboots of the OCA Router and File Server SSC laptops.

The crew performed their regular 2-hr physical exercise on the CEVIS cycle ergometer (FE-2, FE-4, FE-5), TVIS treadmill (CDR, FE-1, FE-3), ARED advanced resistive exerciser (CDR, FE-1, FE-2, FE-4, FE-5), and VELO cycle ergometer with bungee cord load trainer (FE-3).

Afterwards, the CDR transferred the exercise data files to the MEC for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).

At ~11:05am EDT Frank powered up the SM’s amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply) and at 11:10am conducted a ham radio session with Belayr Pathfinders at Dartmouth/Halifax, NS, Canada.

Weekly Science Update (Expedition Twenty-One — Week 2)

3-D SPACE: No report.

AgCam (Agricultural Camera): No report.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): “Ciao a tutti from the science team in Rome and Hello to all from the engineering teams in Rome and at JSC. Thank you for the attention and care in moving the ALTEA Dosi Data Acquisition Unit (DAU). We are continuing to receive good data. All systems are nominal. Grazie!”

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): No report.

BIOLAB (ESA): No report.

BIORHYTHMS (JAXA, Biological Rhythms): On 10/20-21 Jeff’s first session was completed.

BISE (CSA, Bodies in the Space Environment): No report.

BISPHOSPHONATES: “Bob, thanks for completing your pill ingestion. Your next session is scheduled for 10/26. Jeff, thanks for completing your pill ingestion. Your next session is scheduled for 10/25.”

CARD (Long Term Microgravity: Model for Investigating Mechanisms of Heart Disease, ESA): No report.

CARDIOCOG-2: Complete.

CB (JAXA Clean Bench): No report.

CBEF (JAXA Cell Biology Experiment Facility)/SPACE SEED: “The 63-day long experiment began on 9/10. The seedlings in the PEU units are growing well. The first harvesting was conducted on 10/13. CBEF manual ventilation was conducted to lower the humidity inside incubators. Now humidity inside and outside (cabin) is almost same level. Midterm harvested samples in KFT was transferred from 4 degC to -95 degC as scheduled. Final harvesting is now scheduled on 11/10.”

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): No report.

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Complete.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack), MDCA/Flex: No report.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial 2 (JAXA): Completed.

Commercial 3 (JAXA): Completed.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DomeGene (JAXA): Complete.

DOSIS (ESA): No report.

EarthKAM (Earth Knowledge Acquired by Middle School Students): EarthKAM successfully conducted 5-days of image gathering for 122 Education Groups (Elementary through Collegiate levels) for 9761 students. 97 of the groups are from US Schools, 10 from Australia, the remainder from Spain: 2, Germany: 2, Japan, Dominican Republic, France, Lebanon, United Kingdom, Finland, India, Denmark, Netherlands, Romania, Guatemala. Over 750 images have been downlinked with two more days of operations still to follow. “Image quality is superb!”

EDR (European Drawer Rack, ESA): The rack is continuously active in support of the Protein Crystallization Diagnostic Facility (PCDF) experiment. EDR is providing power/data and temperature control (via cooling loop) to PCDF.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

ENose (Electronic Nose): No report.

EPM (European Physiology Module): No report.

EPO (Educational Payload Operations, NASA): No report.

EPO LES-3 (ESA): No report.

EPO 3-min Video (JAXA): No report.

EPO J-Astro Report (JAXA): Ongoing.

EPO Dewey’s Forest (JAXA): The sample was transferred to MELFI.

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): No report.

EPO Moon Score (JAXA): No report.

EPO Try Zero-G (JAXA): “No report.

EPO Kibo Kids Tour (JAXA): Complete.

EPO Spiral Top (JAXA): No report.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): With landing of 17A on 9/11, EuTEF platform was returned to the ground.

FACET (JAXA): No report.


FSL (Fluid Science Laboratory): No report.

GEOFLOW: No report.

HDTV System (JAXA): To be launched by HTV1.

Holter ECG (JAXA): No report.

HQPC (JAXA): To be launched by 34P.

HREP (HICO/Hyperspectral Imager for the Coastal Ocean & RAIDS/Remote Atmospheric & Ionospheric Detection System/JAXA): No report.


ICV (Integrated Cardiovascular): “Jeff and Nicole, the entire team enjoyed participating in your first ICV resting echo scan. Nicole did a great job as the operator and Jeff’s heart did just what it was supposed to – beat and provide great images! Since the experiment requires two sessions early in flight (FD14 & FD30) we will do another Echo in a little over a week.”

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.


InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): No report.

IRIS (Image Reversal in Space): No report.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): No report.

MAXI (Monitor of All-sky X-ray Image, JAXA): No report.

MEIS (Marangoni Experiment for ISS) in JAXA FPEF (Fluid Physics Experiment Facility): Postponed because of sample cassette failure.

MDCA/Flex: See under CIR.

MDS (Mice Drawer System): “Good day Nicole, Bob and team. We are confident that the continued nominal function of the MDS facility and the good health of the mice are directly related to your care and attention of this science investigation. This past week was very good with respect to all system parameters and the telemetry concerning the mice. This coming week we expect to remove the muffler for a 24-hr period to study the effects on the humidity within the MDS facility. Thank you very much for everything which you continue to do on behalf of the MDS teams on the ground.”

Microbe-1 (JAXA): No report.

Micro-G Clay (JAXA EPO): Complete.

MMA (JAXA/Microgravity Measurement Apparatus): No report.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): No report.

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC-1 “Pirs”.

MULTIGEN-1: Completed.

NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): No report.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.


PADLES (JAXA, Area PADLES 3; Passive Area Dosimeter for Lifescience Experiment in Space): No report.

PCDF-PU (Protein Crystallization Diagnostic Facility – Process Unit): No report.

PCG (JAXA, Protein Crystal Growth): PCG samples were returned by Soyuz 18S and transported to JAXA.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.


RadGene & LOH (JAXA): Complete.

RadSilk (JAXA): RadSilk experiment has started. Sortie sample (launch control sample) was returned by 17A.

RST/Reaction Self Test (Psychomotor Vigilance Self Test on the ISS): “Jeff, thank you for completing the first Reaction Self Test sleep shift sequence. Nominal Reaction Self Test sessions are scheduled every four days. Thank you for your participation and we look forward to analyzing the data from your sessions.”

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SEDA-AP (Space Environment Data Acquisition Equipment-Attached Payload, JAXA): Started nominal operation.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “All, next week we are targeting your next week of sleep logging. It will be the first week for this increment. Bob, thanks for downloading and initializing all the Actiwatches. All the data looked good except for Mike’s. His Actiwatch stopped taking data 10 days after initialization. HRF does not plan to use this Actiwatch until troubleshooting can be completed.”

SLAMMD (Space Linear Acceleration Mass Measurement Device): No report.

SMILES (JAXA): No report.

SODI/IVIDIL (Selectable Optical Diagnostics Instrument/Influence of Vibration on Diffusion in. Liquids, ESA): SODI-IVIDIL runs continue this week on the first cell array. The science team is very happy with the good results they are seeing and this is allowing them to make some modifications to future runs.

SOLAR (Solar Monitoring Observatory): Current Sun observation window started on 10/15 and is predicted to end on 10/28. Some observation orbits were lost when SOLAR had to be put in safe configuration for the 35P docking and the MSL venting events.

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): No report.

SPICE (Smoke Point In Co-flow Experiment): No report.

SPINAL (Spinal Elongation): No report.

SWAB (Characterization of Microorganisms & Allergens in Spacecraft): “Nicole and Frank, thank you for completing the third of eight SWAB water collections! The expiration dates you noticed on the hardware were deemed unnecessary by the SWAB PI team after the hardware launched. A note on this is in the Big Picture words for SWAB. Big Picture words also provide a suggestion to shake the fixative powder tubes to break up any clumps before beginning water transfer. It might also help to release some powder from the tube after beginning the water transfer but before the water reaches the bottom portion of the bag. These techniques were suggested by the SWAB PI team after feedback from the first two SWAB sessions. The pictures of SWAB activities are great! We appreciate being able to see the operations. We thank you both again for working together to get the sample collected. We look forward to analyzing the samples when they return on STS 129/ULF3. Thank you for your work!”

THERMOLAB (ESA): No report.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


VO2max (NASA): “Jeff, we are happy to report that we have received approval to leave most of the Portable PFS hardware deployed in a temporary stow configuration when ops are planned to occur in the near future.”

VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage). No report.

YEAST B (ESA): No report.

CEO (Crew Earth Observations): Through 10/21, the ground has received a total of 5,354 frames of ISS-21 CEO imagery for review and cataloguing (including over 4,000 frames in the past week alone). “We are pleased to confirm your acquisition of imagery for the following CEO targets: Lake Eyre, Australia – 2 excellent frames, the first in over a year – good documentation of the slow refilling of this large, salt lake; Hyderabad, India – several useful frames of portions of this urban area; and Lake Poopo, Bolivia – a single context frame. Your recent, detailed view of the Great Bitter Lake near the middle of Egypt’s Suez Canal will be published on NASA/GSFC’s Earth Observatory website this weekend. It nicely documents a key location on this vital commercial waterway. Nice shot! Please feel free to advise us if we can be of any further assistance in helping you locate specific target requests, or, if you need any additional feedback on your imagery. Thanks for your support of our payload.”

CEO (Crew Earth Observation) photo targets uplinked for today were Popocatepetl Volcano, Mexico (looking right of track for this large volcano located 70 km to the SE of Mexico City. Three major explosive eruptions have occurred in the very recent geologic past, producing pyroclastic flows and lahars [mud flows] that affected the basins surrounding the volcano. Mapping frames of the volcano and flanks are requested to capture current summit glacier extent and cone geomorphology. Gas and steam emissions can at times be observed emanating from the volcano), and Cape Tres Montes, Chile (HMS Beagle Site: Cape Tres Montes was as far south in South America as the Beagle would travel. In his letters, Darwin notes the numerous forested islands and mountains. Looking left of track, mapping pass along the cape).

ISS Orbit (as of this morning, 8:31am EDT [= epoch])
Mean altitude – 343.6 km
Apogee height – 348.0 km
Perigee height – 339.3 km
Period — 91.41 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0006431
Solar Beta Angle — 10.1 deg (magnitude increasing)
Orbits per 24-hr. day — 15.75
Mean altitude loss in the last 24 hours — 223 m
Revolutions since FGB/Zarya launch (Nov. 98) — 62632

Significant Events Ahead (all dates Eastern Time, some changes possible!):
10/27/09 — Ares I-X Flight Test (8:00am EDT)
10/29/09 — HTV1 hatch closing
10/30/09 — HTV1 unberthing (12:05pm EDT)
11/01/09 — Daylight Time ends/Standard Time begins
11/04/09 — HTV1 reentry (destructive)
11/10/09 — 5R/MRM-2 (Russian Mini Research Module 2) launch on Soyuz-U
11/12/09 — 5R/MRM-2 docking (SM zenith)
NET 11/16/09 — STS-129/Atlantis/ULF3 launch (ELC1, ELC2) 2:28pm EST — not earlier than
12/01/09 – Soyuz TMA-15/19S undock
12/01-12/23 —> two-member crew
12/21/09 — Soyuz TMA-17/21S launch — O. Kotov/S. Noguchi/T.J. Creamer
12/23/09 — Soyuz TMA-17/21S (FGB nadir)
01/20/10 — Soyuz TMA-16/20S relocation (from SM aft to MRM-2)
02/03/10 — Progress M-04M/36P launch
02/04/10 — STS-130/Endeavour/20A – Node-3 + Cupola
02/05/10 — Progress M-04M/36P docking
03/18/10 — Soyuz TMA-16/20S undock/landing
03/18/10 — STS-131/Discovery/19A – MPLM(P), LMC
04/02/10 — Soyuz TMA-18/22S launch
04/27/10 — Progress M-03M/35P undock
04/28/10 — Progress M-05M/37P launch
04/30/10 — Progress M-05M/37P docking
05/14/10 — STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1
05/29/10 — Progress M-04M/36P undock
05/30/10 — Soyuz TMA-19/23S launch
06/30/10 — Progress M-06M/38P launch
07/02/10 — Progress M-06M/38P docking
07/26/10 — Progress M-05M/37P undock
07/27/10 — Progress M-07M/39P launch
07/29/10 — Progress M-07M/39P docking
07/29/10 — STS-134/Endeavour (ULF6 – ELC3, AMS-02)
08/30/10 — Progress M-06M/38P undock
08/31/10 — Progress M-08M/40P launch
09/02/10 — Progress M-08M/40P docking
09/16/10 — STS-133/Discovery (ULF5 – ELC4, PLM)
09/18/10 — STS-133/Discovery (ULF5 – ELC4, PLM) docking
09/22/10 — STS-133/Discovery (ULF5 – ELC4, PLM) undock
09/30/10 — Soyuz TMA-20/24S launch
10/26/10 — Progress M-07M/39P undock
10/27/10 — Progress M-09M/41P launch
10/29/10 — Progress M-09M/41P docking
11/30/10 — ATV2 launch– Ariane 5 (ESA)
11/30/10 — Soyuz TMA-21/25S launch
12/15/10 — Progress M-08M/40P undock
12/17/10 — ATV2 docking
02/08/11 — Progress M-09M/41P undock
02/09/11 — Progress M-10M/42P launch
02/11/11 — Progress M-10M/42P docking
03/30/11 — Soyuz TMA-22/26S launch
xx/xx/11 – Progress M-11M/43P launch
05/30/11 — Soyuz TMA-23/27S launch
12/??/11 — 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton

SpaceRef staff editor.